OpenStudy (anonymous):

how do i put this in y=mx+b form; 2x+4y=-4

OpenStudy (anonymous):

y is 4y x is 2x m is -4 so 4y = -4 (2x) + 0 >>> -8x + 0 divide by 4 we get y = -4x

OpenStudy (anonymous):

^ Good try, but the answer is: Solve for y: \[2x + 4y = 4\] \[4y = -2x +4\] \[y = -x/2+1\] Therefore, m = -1/2 & b = 4

OpenStudy (anonymous):

aha i see ! i got the answer incorrectly ><" sorry